OpenStudy (anonymous):

If the density of 1.77 grams of gold is 19.3 g/ml what's the volume?

OpenStudy (whpalmer4):

V = 1.77 g / 19.3 g/ml =

OpenStudy (souvik):

D=M/V or V=M/D where V=volume M=mass D=density